# What Can Large Language Models Do for Sustainable Food?

### v1

@@ Abstract
Food systems are responsible for a third of human-caused greenhouse gas emissions. We investigate what Large Language Models (LLMs) can contribute to reducing the environmental impacts of food production. We define a typology of design and prediction tasks based on the sustainable food literature and collaboration with domain experts, and evaluate six LLMs on four tasks in our typology. For example, for a sustainable protein design task, food science experts estimated that collaboration with an LLM can reduce time spent by 45% on average, compared to 22% for collaboration with another expert human food scientist. However, for a sustainable menu design task, LLMs produce suboptimal solutions when instructed to consider both human satisfaction and climate impacts. We propose a general framework for integrating LLMs with combinatorial optimization to improve reasoning capabilities. Our approach decreases emissions of food choices by 79% in a hypothetical restaurant while maintaining participants’ satisfaction with their set of choices. Our results demonstrate LLMs’ potential, supported by optimization techniques, to accelerate sustainable food development and adoption.
@@ 1 Introduction
Global food production practices are a leading contributor to climate change, deforestation, water pollution, biodiversity loss, antibiotic resistance, and zoonotic disease transmission (Allan et al., 2023; Hopwood et al., 2020; Scherer et al., 2019). Given the disproportionate harms of industrial animal agriculture in particular, which accounts for 57% of food systems emissions, a number of companies have formed to develop more sustainable protein sources (Xu et al., 2021; Mylan et al., 2023). Additionally, programs such as the Menus of Change have developed to promote sustainable choices in foodservice operations (Menus of Change, 2024). These directions require knowledge of both food science and human preferences.
@@ 2 Related Work
Large Language Model Evaluations. Rigorous evaluation of LLMs has clarified their strengths and weaknesses. Guha et al. (2024) identified six types of legal reasoning tasks and evaluated 20 models on them. In chemistry, Guo et al. (2023) identified three types of capabi, established a benchmark of eight tasks, and evaluated five LLMs. However, no work, to our knowledge, has formalized key tasks and evaluated LLMs in the sustainable food domain.
@@ 5.2 Menu Design (Traditional Foods)
As a reference menu, we use a 36-item menu from Banerjee et al. (2023), created using Deliveroo’s and Just Eat’s top 100 items ordered in the United Kingdom in 2019. Each LLM was instructed to reduce emissions of food choices by 75% while maintaining satisfaction, cost, nutrition, preparation time, and animal welfare, and using only the same ingredients as the original menu other than standard vegetarian protein sources (most of which were already in the original menu, other than eggs and chickpeas). Adherence to the ingredient availability constraint, as well as meeting the constraint of generating 36 properly formatted recipes, was checked, and the LLM was given up to five chances to correct itself. GPT-3.5 Turbo and Llama 3.1 70b Instruct did not meet the formatting or ingredient constraints.
@@ 5.3 Sensory Profile Prediction (Sustainable Protein)
Results. Accuracies are shown in Table 2. Accuracies significant according to a chi-squared test with Bonferroni correction for multiple testing are bolded (VanderWeele & Mathur, 2019). Performance is generally not meaningfully superior to the baseline. Accuracy for the best (across all dimensions) LLM, o1-preview, improves in the highest quartile of the ground truth difference between the recipes, e.g. to 81% and 75% for overall satisfaction and purchase intent. However, it is still not meaningfully superior to the simple baseline, which achieved 86% and 70% accuracy on those dimensions in this subset of pairs (Appendix Table 16).
@@ 5.4 Recipe Preference Prediction (Traditional Foods)
Methods. We created a set of 500 meal recipe pairs with similar ingredients but significantly different average ratings ($p<0.05$ in a $t$-test), ensuring that one recipe is consistently preferred by users. To identify recipe pairs with the highest similarity, the overlap of ingredients was calculated for each pair using the ratio of shared ingredients to the total number of unique ingredients. We addressed the possibility of overlap with LLM training data by curating an additional set of 58 recipe pairs from epicurious.com. These recipes were posted behind a paywall after the release dates of all the models we tested. Our primary metric was accuracy—whether the models correctly identified the preferred recipe—compared to a random baseline. We tested statistical significance using a chi-squared test with Bonferroni correction when evaluating on multiple conditions. The order in which the recipes were shown was randomized.
Results. Claude 3.5 Sonnet, GPT-4o, and o1-preview outperform the random baseline (Table 2). The highest accuracy achieved is 64% (GPT-4o), with accuracy improving to 85% in the highest quartile of the ground truth review score gap (Appendix Table 17). Moreover, performance varies widely by comparison type (Appendix Table 18). Across all the models, accuracy is the highest when two non-vegetarian meals are compared, followed by vegetarian vs. non-vegetarian comparisons, with the worst accuracy observed in vegetarian vs. vegetarian meal comparisons. This aligns with existing literature showing that minority preferences and opinions are modeled less accurately (Santurkar et al., 2023). Within the vegetarian vs. non-vegetarian meal comparisons, accuracy is lower when vegetarian options are rated higher, reflecting LLMs’ omnivore bias, as LLMs more frequently prefer non-vegetarian options compared to ground truth human preferences. This finding is consistent with the literature on bias amplification in the food domain (Luo et al., 2024) and points to the need to improve preference prediction performance in order to develop tools that can better support the development of sustainable food. Finally, similar pairwise accuracy across models on post-cut-off data suggests that performance cannot be fully attributed to memorization (Appendix Table 19).
@@ 6 Augmenting LLMs with Combinatorial Optimization
We propose to address these challenges by combining LLMs’ background knowledge and ability to model human preferences, demonstrated in Horton (2023)Park et al. (2022) and in our own experiments, particularly for coarse-grained prediction (Tables 2,10,16,17), with mathematical optimization. While LLMs alone struggle with optimization, and traditional optimization methods do not have access to LLMs’ background knowledge, combining them allows for leveraging the strengths of both approaches.
@@ 6.1 LLM-Guided Combinatorial Optimization
Let $x^{\ast}$ be the optimal solution with respect to actual preferences $p(u_{i})$, and $\hat{x}^{\ast}$ be the optimal solution with respect to estimated preferences $\hat{p}(u_{i})$. Let $p(u_{i}),\hat{p}(u_{i})\in[0,1]$, $d(x)\in[0,D]$. If $|\hat{p}(u_{i})-p(u_{i})|\leq\epsilon$ for all $u_{i}\in U$, then $|f(\hat{x}^{\ast})-f(x^{\ast})|\leq 2K\epsilon+2\lambda D$.
@@ 6.2 Constrained Menu Design Formulation
Now we show how to implement this framework for the setting of constrained menu design. First, we prompt the LLM to generate a set $C$ of plant-based candidate recipes, though this could be modified to allow for also generating animal-based recipes at this stage. As in Section 5.2, the LLM is also instructed to not worsen cost, nutrition, or preparation time. We then combine this set with the recipes in the original menu $O$, where $K=|O|$, to form the ground set $U=C\cup O$. We then estimate $p(u_{i})\>\forall u_{i}\in U$ by prompting an LLM to rate each recipe in the ground set by expected preferences of the population of interest (here, American omnivores). Then, integer quadratic programming (IQP) is applied to select a subset of recipes from $U$ for the menu.
@@
We set $d(x)$ to be negative pairwise similarity between the recipes in the selected subset. This yields an IQP since all other terms are linear in $x$. $C_{i}$ controls the reduction in emissions or animal welfare, and is set to 0.25 for emissions and 1.0 for animal welfare, corresponding to reducing emissions by at least 75% while maintaining or improving upon animal welfare. We set $K=36$, to match the length of the original menu from Banerjee et al. (2023). We instruct the LLM to generate 20 additional recipes, reflecting that a restaurant may not be able to completely revise their full menu. Thus, $N=56$. Finally, we set $\lambda=100$. For computing pairwise recipe similarity we use the Ratcliff/Obershelp sequence matching algorithm as implemented in Python’s difflib (Ratcliff et al., 1988). The recipes were ordered first by whether they were LLM generated, and second by predicted rating. We included a baseline in which the original menu was simply re-ordered so that vegetarian items were first, to account for the possibility that benefits of the LLM+IQP approach were simply due to placing the LLM generated (and therefore plant-based) recipes first.
@@
We then evaluated our approach via two human subjects experiments (total $n=552$) in which participants recruited via Prolific, constrained to be in the United States and fluent English speakers, were instructed to make a selection from a menu. In the first experiment, the menu was randomly assigned from one of the five LLMs (augmented with IQP) and three baselines (original menu, original menu with beef removed, and a vegetarian subset of the original menu). [^3] Removal of beef is a common strategy for reducing climate impacts (Grummon et al., 2023). In the second experiment, conducted the following day, the menu was randomly assigned from one of three additional baselines (o1-preview on its own, without IQP; o1-preview+IQP, but with the o1-preview generated descriptions replaced with a simple list of ingredients; and placing the vegetarian items first in the original menu). The participants were paid \$16 per hour. The evaluation metrics consisted of satisfaction metrics (satisfaction with their set of choices, likelihood of visiting, and likelihood of recommending the restaurant to a friend), emissions of food choices, and animal welfare impacts of food choices.
@@ 6.3 Results
The results for the best LLM, o1-preview, are shown in Figure 2, with results for other LLMs and baselines in Appendix E.2. When controlling for dietary preference, age, gender, and race, we estimate that the o1-preview+IQP approach reduces emissions by 79% while maintaining patron satisfaction with their set of choices ($p$=1.37e-18). We obtain similar results when removing vegan and vegetarian participants (Appendix E.2). We observe a small (5%) reduction in the likelihood of recommending the restaurant to a friend, and a small increase in animal usage (3.5%), but these were not statistically significant. The original, baseline, and generated menus are shown in Appendix F. We also evaluated the accuracy of the LLM’s predicted preference ratings, in the same pairwise ranking fashion as the sensory profile and recipe preference prediction tasks, with the best LLM, Claude 3.5 Sonnet, achieving 73% accuracy (Table 2). This increases to 85% in the highest quartile of the preference gap (Appendix Table 10). We note that, for the menu design task, LLM predictions were based on the recipe title alone (not including the full description), which could explain the discrepancy between item-level prediction performance and menu-level performance for o1-preview. When o1-preview’s descriptions were replaced with a simple list of ingredients (in one of our ablations), its accuracy was 70%, and 84% in the highest quartile of ground truth preference gaps (Appendix Table 10).
@@ 7 Discussion
We evaluated LLMs on three types of preference prediction. The first is sensory profile prediction, in which we compared LLM predictions to evaluations from an actual taste tasting. The second is the Food.com recipe rating task, where we compared LLM predictions to online recipe ratings. The third is recipe rating in the context of the menu design experiment, where the LLM’s recipe-level predictions are compared against the order frequency for the recipes that were selected for the final menu. For the holistic measures of overall satisfaction and purchase intent in the sensory profile prediction task, no LLMs outperform a random baseline, though some do on specific dimensions such as greasiness or saltiness. For online recipe rating, three of six LLMs outperform a random baseline. We note that all evaluated LLMs outperform a random baseline in context of the menu design experiment, where no actual tasting occurs and the orders are purely based on the text of the recipe (Table 2). Additionally, we studied how performance varies with the magnitude of the ground truth preference gap. When restricting the evaluated recipe pairs to be in the top quartile of preference gaps, LLM performance improves, with the best LLMs achieving 81%, 86%, and 85% accuracy respectively in the three prediction tasks (Appendix Tables 10,16,17). Thus, our analysis suggests that LLMs can be useful for coarse-grained preference prediction tasks, but that further work is needed for fine-grained prediction, particularly for the mapping from recipes or product formulations to actual taste. Additionally, consistent with past research on LLMs’ modeling of minority preferences (Santurkar et al., 2023), our results showed that LLMs model human preferences related to plant-based food less accurately. Thus, regular human feedback, e.g. taste testings, from diverse populations remains critical. Overall, these results suggest that LLMs can be useful in experimental design, generation of sustainable recipes (as evaluated in the menu design task), and coarse-grained preference prediction.
@@
In the menu design task, we consider ethical aspects such as the trade-offs among patron satisfaction, sustainability, and animal welfare, and show that our approach can improve or maintain all dimensions. Designing menus to shift the distribution of patron choices may raise questions of respect for individual autonomy. We thus use three different metrics of patron satisfaction, and show that all are maintained with the revised menu (Appendix Figures 14,15,16).
@@ Acknowledgments
This work is supported by the Stanford Plant-Based Diet Initiative (PBDI). Anna Thomas is supported by National Institutes of Health Grant R01LM013866. Kristina Gligorić is supported by the Swiss National Science Foundation (Grant P500PT-211127). We thank Caroline Cotto and Max Elder for access to the NECTAR dataset, and Moses Charikar, Ali Teshnizi, and Madeleine Udell for helpful discussions. We thank Yifan Mai and the Center for Research on Foundation Models (CRFM) for their support.
@@
Let $p(u_{i}),\hat{p}(u_{i})\in[0,1]$, and $d(x)\in[0,D]$. If $|\hat{p}(u_{i})-p(u_{i})|\leq\epsilon$ for all $u_{i}\in U$, then $|f(\hat{x}^{\ast})-f(x^{\ast})|\leq 2K\epsilon+2\lambda D$.
@@
First, note that $\forall x,|f(x)-\hat{f}(x)|\leq K\epsilon+\lambda D$. Then, combining this with the optimality of $x^{\ast}$ and $\hat{x}^{\ast}$, $f(\hat{x}^{\ast})\geq\hat{f}(\hat{x}^{\ast})-(K\epsilon+\lambda D)\geq\hat{f}(x^{\ast})-(K\epsilon+\lambda D)\geq f(x^{\ast})-2K\epsilon-2\lambda D$. ∎

### v2

@@ Abstract
Food systems are responsible for a third of human-caused greenhouse gas emissions. We investigate what Large Language Models (LLMs) can contribute to reducing the environmental impacts of food production. We define a typology of design and prediction tasks based on the sustainable food literature and collaboration with domain experts, and evaluate six LLMs on four tasks in our typology. For example, for a sustainable protein design task, food science experts estimated that collaboration with an LLM can reduce time spent by 45% on average, compared to 22% for collaboration with another expert human food scientist. However, for a sustainable menu design task, LLMs produce suboptimal solutions when instructed to consider both human satisfaction and climate impacts. We propose a general framework for integrating LLMs with combinatorial optimization to improve reasoning capabilities. Our approach decreases emissions of food choices by 79% in a hypothetical restaurant while maintaining participants’ satisfaction with their set of choices. Our results demonstrate LLMs’ potential, supported by optimization techniques, to accelerate sustainable food development and adoption. Our code is available at https://github.com/thomasat/llms-sustainable-food.
@@ 1 Introduction
Global food production practices are a leading contributor to climate change, deforestation, water pollution, biodiversity loss, antibiotic resistance, and zoonotic disease transmission (Allan et al., 2023; Hopwood et al., 2020; Scherer et al., 2019). Given the disproportionate harms of animal agriculture in particular, which accounts for 57% of food systems emissions, a number of companies have formed to develop more sustainable protein sources (Xu et al., 2021; Mylan et al., 2023). Additionally, programs such as the Menus of Change have developed to promote sustainable choices in foodservice operations (Menus of Change, 2024). These directions require knowledge of both food science and human preferences.
@@ 2 Related Work
Large Language Model Evaluations. Rigorous evaluation of LLMs has clarified their strengths and weaknesses. Guha et al. (2024) identified six types of legal reasoning tasks and evaluated 20 models on them. In chemistry, Guo et al. (2023) identified three types of capabilities, established a benchmark of eight tasks, and evaluated five LLMs. However, no work, to our knowledge, has formalized key tasks and evaluated LLMs in the sustainable food domain.
@@ 5.2 Menu Design (Traditional Foods)
As a reference menu, we use a 36-item menu from Banerjee et al. (2023), created using Deliveroo’s and Just Eat’s top 100 items ordered in the United Kingdom in 2019. Each item consists of a title and an appealing description, e.g. “Chicken Curry Ramen. Japanese fried chicken and noodles in a delicious curry broth.” Each LLM was instructed to reduce emissions of food choices by 75% while maintaining satisfaction, cost, nutrition, preparation time, and animal welfare, and using only the same ingredients as the original menu other than standard vegetarian protein sources (most of which were already in the original menu, other than eggs and chickpeas). Adherence to the ingredient availability constraint, as well as meeting the constraint of generating 36 properly formatted recipes, was checked, and the LLM was given up to five chances to correct itself. GPT-3.5 Turbo and Llama 3.1 70b Instruct did not meet the formatting or ingredient constraints.
@@ 5.3 Sensory Profile Prediction (Sustainable Protein)
Results. Accuracies are shown in Table 2. Accuracies significant according to a chi-squared test with Bonferroni correction for multiple testing are bolded (VanderWeele & Mathur, 2019). Performance is generally not meaningfully superior to the baseline. Accuracy for the best (across all dimensions) LLM, o1-preview, improves in the highest quartile of the ground truth difference between the recipes, e.g. to 81% and 75% for overall satisfaction and purchase intent. However, it is still not meaningfully superior to the simple baseline, which achieved 86% and 70% accuracy on those dimensions in this subset of pairs (Appendix Table 18). An expert human baseline for the “Overall Satisfaction” dimension was collected from a collaborator with deep domain expertise (11 years of food science experience, including co-founding two plant-based meat companies), who is also a co-author on this paper. The results are shown in Appendix Table 14.
@@ 5.4 Recipe Preference Prediction (Traditional Foods)
Methods. We created a set of 500 recipe pairs with similar ingredients but significantly different average ratings ($p<0.05$ in a $t$-test), ensuring that one recipe is consistently preferred by users. Examples pairs are shown in Appendix C.2. To identify recipe pairs with the highest similarity, the overlap of ingredients was calculated for each pair using the ratio of shared ingredients to the total number of unique ingredients. We addressed the possibility of overlap with LLM training data by curating an additional set of 58 recipe pairs from epicurious.com. These recipes were posted behind a paywall after the release dates of all the models we tested. Our primary metric was accuracy—whether the models correctly identified the preferred recipe—compared to a random baseline. We tested statistical significance using a chi-squared test with Bonferroni correction when evaluating on multiple conditions. The order in which the recipes were shown was randomized.
Results. Claude 3.5 Sonnet, GPT-4o, and o1-preview outperform the random baseline (Table 2). The highest accuracy achieved is 64% (GPT-4o), with accuracy improving to 85% in the highest quartile of the ground truth review score gap, i.e., pairs of recipes for which people’s preferences were most different (Appendix Table 19). Moreover, performance varies widely by comparison type (Appendix Table 20). Across all the models, accuracy is the highest when two non-vegetarian meals are compared, followed by vegetarian vs. non-vegetarian comparisons, with the worst accuracy observed in vegetarian vs. vegetarian meal comparisons. This aligns with existing literature showing that minority preferences and opinions are modeled less accurately (Santurkar et al., 2023). Within the vegetarian vs. non-vegetarian meal comparisons, accuracy is lower when vegetarian options are rated higher, reflecting LLMs’ omnivore bias, as LLMs more frequently prefer non-vegetarian options compared to ground truth human preferences. This finding is consistent with the literature on bias amplification in the food domain (Luo et al., 2024) and points to the need to improve preference prediction performance in order to develop tools that can better support the development of sustainable food. Finally, similar pairwise accuracy across models on post-cut-off data suggests that performance cannot be fully attributed to memorization (Appendix Table 21).
@@ 6 Augmenting LLMs with Combinatorial Optimization
We propose to address these challenges by combining LLMs’ background knowledge and ability to model human preferences, demonstrated in Horton (2023)Park et al. (2022) and in our own experiments, particularly for coarse-grained prediction (Tables 2,11,18,19), with mathematical optimization. While LLMs alone struggle with optimization, and traditional optimization methods do not have access to LLMs’ background knowledge, combining them allows for leveraging the strengths of both approaches.
@@ 6.1 LLM-Guided Combinatorial Optimization
Let $x^{\ast}$ be the optimal solution with respect to actual preferences $p(u_{i})$, and $\hat{x}^{\ast}$ be the optimal solution with respect to estimated preferences $\hat{p}(u_{i})$. Let $p(u_{i}),\hat{p}(u_{i})\in[0,1]$, $d(x)\in[0,D]$. If $|\hat{p}(u_{i})-p(u_{i})|\leq\epsilon$ for all $u_{i}\in U$, then $|f(\hat{x}^{\ast})-f(x^{\ast})|\leq 2K\epsilon$.
@@ 6.2 Constrained Menu Design Formulation
Now we show how to implement this framework for the setting of constrained menu design. First, we prompt the LLM to generate a set $C$ of plant-based candidate recipes, though this could be modified to allow for also generating animal-based recipes at this stage. As in Section 5.2, the LLM is also instructed to not worsen cost, nutrition, or preparation time. We then combine this set with the recipes in the original menu $O$, where $K=|O|$, to form the ground set $U=C\cup O$. We then estimate $p(u_{i})\>\forall u_{i}\in U$ by prompting an LLM to rate each recipe title in the ground set by expected preferences of the population of interest (here, American omnivores). Then, integer quadratic programming (IQP) is applied to select a subset of recipes from $U$ for the menu.
@@
We set $d(x)$ to be negative pairwise similarity between the recipes in the selected subset. This yields an IQP since all other terms are linear in $x$. $C_{i}$ controls the reduction in emissions or animal welfare, and is set to 0.25 for emissions and 1.0 for animal welfare, corresponding to reducing emissions by at least 75% while maintaining or improving upon animal welfare. We set $K=36$, to match the length of the original menu from Banerjee et al. (2023). We prompt the LLM to generate 20 additional recipes. Thus, $N=56$. Finally, we set $\lambda=100$. For computing pairwise recipe similarity we use the Ratcliff/Obershelp sequence matching algorithm as implemented in Python’s difflib (Ratcliff et al., 1988). The recipes were ordered first by whether they were LLM generated, and second by predicted rating. We included a baseline in which the original menu was simply re-ordered so that vegetarian items were first, to account for the possibility that benefits of the LLM+IQP approach were simply due to placing the LLM generated (and therefore plant-based) recipes first.
@@
We then evaluated our approach via two human subjects experiments (total $n=800$) in which participants recruited via Prolific, constrained to be in the United States and fluent English speakers, were instructed to make a selection from a menu. In the first experiment, the menu was randomly assigned from one of the five LLMs (augmented with IQP) and three baselines (original menu, original menu with beef removed, and a vegetarian subset of the original menu). [^3] Removal of beef is a common strategy for reducing climate impacts (Grummon et al., 2023). In the second experiment, conducted the following day, the menu was randomly assigned from one of three additional baselines (o1-preview on its own, without IQP; o1-preview+IQP, but with the o1-preview generated descriptions replaced with a simple list of ingredients; and placing the vegetarian items first in the original menu). The participants were paid \$16 per hour. The evaluation metrics consisted of satisfaction metrics (satisfaction with their set of choices, likelihood of visiting, and likelihood of recommending the restaurant to a friend), emissions of food choices, and animal welfare impacts of food choices.
@@ 6.3 Results
The results for the best LLM, o1-preview, are shown in Figure 2, with results for other LLMs and baselines in Appendix E.2. When controlling for dietary preference, age, gender, and race, we estimate that the o1-preview+IQP approach reduces emissions by 79% while maintaining patron satisfaction with their set of choices ($p$=1.37e-18). We obtain similar results when removing vegan and vegetarian participants (Appendix E.2). We observe a small (5%) reduction in the likelihood of recommending the restaurant to a friend, and a small increase in animal usage (3.5%), but these were not statistically significant. The original, baseline, and generated menus are shown in Appendix F. We also evaluated the accuracy of the LLM’s predicted preference ratings, in the same pairwise ranking fashion as the sensory profile and recipe preference prediction tasks, with the best LLM, Claude 3.5 Sonnet, achieving 73% accuracy (Table 2). This increases to 85% in the highest quartile of the preference gap (Appendix Table 11). An expert human baseline was collected from a collaborator with deep domain expertise (32 years of culinary experience), who is also a co-author on this paper. The results are shown in Appendix Table 13.
@@ 7 Discussion
We evaluated LLMs on three types of preference prediction. The first is sensory profile prediction, in which we compared LLM predictions to evaluations from an actual taste tasting. The second is the Food.com recipe rating task, where we compared LLM predictions to online recipe ratings. The third is recipe rating in the context of the menu design experiment, where the LLM’s recipe-level predictions are compared against the order frequency - the frequency with which people chose the dishes in our human subjects experiment - for the recipes that were selected for the final menu. For the holistic measures of overall satisfaction and purchase intent in the sensory profile prediction task, no LLMs outperform a random baseline, though some do on specific dimensions such as greasiness or saltiness. For online recipe rating, three of six LLMs outperform a random baseline. All evaluated LLMs outperform a random baseline in context of the menu design experiment, where no actual tasting occurs and the orders are purely based on the text of the recipe (Table 2). Additionally, we studied how performance varies with the magnitude of the ground truth preference gap. When restricting the evaluated recipe pairs to be in the top quartile of preference gaps, LLM performance improves, with the best LLMs achieving 81%, 86%, and 85% accuracy respectively in the three prediction tasks (Appendix Tables 11,18,19). Thus, our analysis suggests that LLMs can be useful for coarse-grained preference prediction tasks, but that further work is needed for fine-grained prediction, particularly for the mapping from recipes or product formulations to actual taste. Additionally, consistent with past research on LLMs’ modeling of minority preferences (Santurkar et al., 2023), our results showed that LLMs model human preferences related to plant-based food less accurately. Thus, regular human feedback, e.g. taste testings, from diverse populations remains critical. Overall, these results suggest that LLMs can be useful in experimental design, generation of sustainable recipes (as evaluated in the menu design task), and coarse-grained preference prediction.
@@
In the menu design task, we consider ethical aspects such as the trade-offs among patron satisfaction, sustainability, and animal welfare, and show that our approach can improve or maintain all dimensions. Designing menus to shift the distribution of patron choices may raise questions of respect for individual autonomy. We thus use three different metrics of patron satisfaction, and show that all are maintained with the revised menu (Appendix Figures 17,18,19).
@@ Acknowledgments
This work is supported by the Stanford Plant-Based Diet Initiative (PBDI). Anna Thomas is supported by National Institutes of Health Grant R01LM013866. Kristina Gligorić is supported by the Swiss National Science Foundation (Grant P500PT-211127). We thank Caroline Cotto and Max Elder for access to the NECTAR dataset, and Moses Charikar, Ali Teshnizi, Madeleine Udell, and Katie Yoon for helpful discussions. We thank Yifan Mai and the Center for Research on Foundation Models (CRFM) for their support.
@@
Let $p(u_{i}),\hat{p}(u_{i})\in[0,1]$, and $d(x)\in[0,D]$. If $|\hat{p}(u_{i})-p(u_{i})|\leq\epsilon$ for all $u_{i}\in U$, then $|f(\hat{x}^{\ast})-f(x^{\ast})|\leq 2K\epsilon$.
@@
First, note that $\forall x,|f(x)-\hat{f}(x)|\leq K\epsilon$. Then, combining this with the optimality of $x^{\ast}$ and $\hat{x}^{\ast}$, $f(\hat{x}^{\ast})\geq\hat{f}(\hat{x}^{\ast})-K\epsilon\geq\hat{f}(x^{\ast})-K\epsilon\geq f(x^{\ast})-2K\epsilon$. ∎